# Supervised Fine-tuning for Cross-Lingual Detoxification

### v1

@@ Abstract
As large language models (LLMs) become increasingly prevalent in global applications, ensuring that they are toxicity-free across diverse linguistic contexts remains a critical challenge. We explore “Cross-lingual Detoxification”, a cross-lingual paradigm that mitigates toxicity, enabling detoxification capabilities to transfer between high and low-resource languages across different script families. We analyze cross-lingual detoxification’s effectiveness through 504 extensive settings to evaluate toxicity reduction in cross-distribution settings with limited data and investigate how mitigation impacts model performance on non-toxic tasks, revealing trade-offs between safety and knowledge preservation. Our code and dataset are publicly available.111 https://github.com/himanshubeniwal/Breaking-mBad
[scale=0.04]pictures/mWildGuard_Logo.pdf
[…]
Supervised Fine-tuning for Cross-Lingual Detoxification
WARNING: The content contains model outputs that are offensive and toxic.
Himanshu Beniwal, Youngwoo Kim, Maarten Sap, Soham Dan∗, Thomas Hartvigsen∗ Indian Institute of Technology Gandhinagar, University of Virginia, Carnegie Mellon University, Allen Institute for Artificial Intelligence, Microsoft Correspondence: himanshubeniwal@iitgn.ac.in
@@ 1 Introduction
| am | ar | de | en | es | hi | ru | uk | zh | Avg | ||
| 15.49 | 19.11 | 20.3 | 30.34 | 22.38 | 27.73 | 19.07 | 22.73 | 22.44 | 22.18 | ||
| am | 4.39 | 5.25 | 0.02 | -0.21 | -0.55 | -0.61 | 0.7 | 0.4 | 1.28 | 1.19 | |
| ar | 2.72 | 2.35 | 1.71 | -1.3 | 6.97 | 8.99 | 2.99 | -1.66 | -5.36 | 1.93 | |
| de | 4.02 | 7.4 | 2.74 | 12.84 | 8.36 | 17.19 | 5.29 | 11.07 | 11.35 | 8.92 | |
| en | -3.5 | -2.25 | -2.41 | 1.77 | -1.41 | 3.32 | 0.08 | -2.51 | -12.87 | -2.20 | |
| es | 1.43 | 10.83 | 7.12 | 16.82 | 8.39 | 16.17 | 5.66 | 13.82 | 7.85 | 9.79 | |
| hi | -1.15 | 0.51 | -8.29 | -16.93 | -8.11 | -6.83 | -12.69 | -7.03 | -14.64 | -8.35 | |
| ru | 4.32 | 3.67 | -1.89 | -1.19 | 0.38 | 0.78 | -0.92 | -2.93 | 2.36 | 0.51 | |
| uk | -1.26 | 3.25 | -2.84 | -9.47 | -2.46 | 11.89 | -13.83 | -6.75 | -0.51 | -2.44 | |
| zh | -1.98 | -2 | -8.08 | -14.65 | -4.97 | -1.11 | -11.31 | -7.52 | -15.33 | -7.44 | |
| - () | Avg | 1.00 | 3.22 | -1.32 | -1.37 | 0.73 | 5.53 | -2.67 | -0.35 | -2.87 |
We investigate Cross-Lingual Detoxification (X-DET), a methodology to detoxify language models in a source language and to evaluate transfer effects across nine target languages. We utilize parallel toxic-neutral pairs to perform the detoxification. We showcase this technique that performs efficiently in cross-lingual settings. Our analysis encompasses 504 experimental configurations, comprising 9 languages (81 language pairs), 4 learning strategies, and 4 mLLMs (details in Section A.2).
Key Findings: Our findings show that: (1) linguistic properties such as morphological complexity and syntactic structures may influence this cross-lingual toxicity transfer in languages with similar scripts and proportion of languages, (2) Models like aya-expanse-8b (Dang et al., 2024) and bloom-7b (Scao et al., 2022), trained on English instances (High-resource langauge), show poor generalization to structurally different languages such as Chinese and Hindi (Figure 2). Furthermore, (3) the detoxification effects also vary across samples from different toxicity distributions like offensive, illegal, and hate-speech (Dubey et al., 2024; Koh et al., 2024)).
@@
Our experiments across 504 configurations show that cross-lingual detoxification significantly outperforms multilingual and proportional fine-tuning approaches.
Cross-distribution detoxification proves effective even with limited parallel data (10%, 20%, and 30% of entire data), achieving effective detoxification without requiring extensive datasets in similar scripts and pretraining language proportion.
@@ 2 Related Work
Early work on identifying and mitigating toxicity in language models focused primarily on English (Gehman et al., 2020; Xu et al., 2021; Leong et al., 2023; Lee et al., 2024). Initial approaches employed supervised fine-tuning with annotated datasets and keyword-based filtering (Pozzobon et al., 2024; Dementieva et al., 2025b), which often degraded model fluency. While subsequent research introduced preference optimization techniques to align models with safety principles (Li et al., 2024), these studies predominantly target high-resource languages, assuming universal transferability of toxicity patterns—an assumption (Moskovskiy et al., 2022; Mukherjee et al., 2023; Wang et al., 2024a; Jain et al., 2024; Jiang and Zubiaga, 2024).
@@ 3 Experiments
Problem Setting
Let be a set of different languages. Each language is associated with a dataset containing pairs of toxic and non-toxic sentences written in language . Detoxification is the task of using toxic sentences from language to update a language model such that it assigns a low probability to toxic sentences across all languages.
Dataset
For our experiments, we utilize the multilingual parallel detoxification dataset: textdetox/multilingual_paradetox222https://huggingface.co/datasets/textdetox/multilingual_paradetox (Bevendorff et al., 2024; Dementieva et al., 2024b, 2025a), which provides parallel toxic and neutral texts across nine333We systematically investigate across the following script families: (1) Latin: German (de), English (en), Spanish (es), (2) Cyrillic: Russian (ru) and Ukrainian (uk), (3) Devnagri: Hindi (hi), (4) Ge’ez: Amharic (am), (5) Arabic: Arabic (ar), (6) Han: Chinese (zh). typologically diverse languages. Each language contains carefully curated parallel samples with toxic content paired with its semantically equivalent neutral (Non-toxic) samples. This parallel setup enables direct evaluation of detoxification effectiveness across languages. More details are in Section §A.1.
Models
@@ 3.1 Learning and Evaluation Paradigms
We evaluate and fine-tune models in four settings:
@@
Cross-lingual Fine-Tuning (-): We finetune each model on 300 neutral-train samples from one source language and evaluate on the test sets across all target languages, yielding 81 language-pair configurations across 9 languages.
@@
Multilingual Fine-Tuning (-): Models are fine-tuned on the aggregated neutral-train samples across all languages (2,700 samples; 300 samples x 9 languages) and evaluated on test sets.
@@ 3.2 Metrics
We prompt the model with the following simple template for all the toxic-train, toxic-test, and neutral-test samples:
Complete the sentence: {SENTENCE}
We evaluate model performance using two standard metrics: (1) toxicity and (2) perplexity. For toxicity detection, we employ Perspective-API,444https://perspectiveapi.com/ a standardized tool supporting 17 languages across high and low-resource categories, to compute the average toxicity score () across languages (Gehman et al., 2020; Jain et al., 2024). We present the toxicity measurements for the zero-shot () baseline and the corresponding mitigation delta scores555The differential mitigation scores () are calculated by computing the arithmetic difference between the toxicity baseline and the respective fine-tuned variants’ toxicity scores (, where ). () for models fine-tuned with -, -, and -. The model’s perplexity is computed using fine-tuned models. More details are provided in Appendix §A.5.
@@ 4 Results and Discussion
We formulate and discuss on the following research questions (RQs):
(RQ1) How well does detoxification transfer across languages? Analysis of the aya-23-8B model in Figure 2 shows superior detoxification in high-resource languages: es (10.6%), de (9%), and en (2.76%), with similar trends in aya-expanse-8B (Table 1). Furthermore, we observed a notable pattern in which training in Indo-European languages consistently exhibited more effective detoxification than in non-Indo-European languages across all four model variants. We attribute this disparity to two primary factors: (1) the proportional representation of languages during the pretraining phase, and (2) the inherent similarities in script among related languages.
@@
(RQ2) Can we effectively mitigate toxicity in cross-lingual settings with limited data? Figure 3 illustrates the variation in toxicity scores across different training data proportions: , , , and (-), where we finetune on the portion of languages and report the AT over a specific language. Notably, am and ar demonstrated improved detoxification performance, aligning with the trends observed in en and es. Our analysis of these languages’ behavior, presented in Figures 7 and 8 (detailed further in Section §A.4), reveals that the fine-tuning causes the embedding representations to converge, suggesting increased similarity in the model’s handling of toxicity across these languages.
@@
(RQ3) How does cross-lingual detoxification impact perplexity? Our perplexity analysis reveals that Indo-European languages, particularly hi, show improved scores (9.01) in aya-expanse-8B’s toxic-train split (Table 15), though both - and - negatively impacted overall perplexity across models (More details in Section §A.5). Embedding similarity analysis before and after detoxification indicates a shift in the relationship between en and am, with their similarity score decreasing to 0.66 in Figures 7 and 8.
@@ 5 Conclusion
Our work explores the challenges of Large Language Models (LLMs) in generating toxic content across different language families, including Indo-European, Non-Indo-European, and Right-to-Left script languages. Given our limited computational resources and the complex nature of our experiments, we had to restrict our analysis to nine languages, four model variants, and four learning strategies. Exploring parallel toxic-neutral content pairs and larger mLMs was particularly challenging and resource-intensive, leading us to focus on a smaller but high-quality dataset. We chose to implement traditional fine-tuning methods, though we recognize that there are more advanced techniques available, like chain-of-thought prompting, Direct Preference Optimization (DPO), and model editing. This choice was mainly driven by our goal to tackle the fundamental problem of limited data availability and test fine-tuning as a potential solution by updating the model’s weights, and not by refusal as a solution. Furthermore, the models are susceptible to jailbreaking, adversarial attacks, and using toxic refusal (ex., “Sorry I cannot respond..”) (Morris et al., 2020). Thus, we prioritized weight updation as a strategy. Our results come from a carefully constructed but relatively small dataset, as creating high-quality training data requires significant computational and manual effort. Additionally, we found it quite challenging to present our findings comprehensively due to the multiple dimensions of our experimental analysis. Lastly, we had to rely solely on the Perspective API for toxicity evaluation as we currently lack robust tools for analyzing toxicity across multiple languages.
@@ A.1 Dataset Split
From the original set, we create our experimental splits by sampling 400 pairs, constructing a training set of 300 parallel pairs (toxic-train and neutral-train) and a test set of 100 pairs (toxic-test and neutral-test). Further, the textdetox/multilingual_paradetox dataset666https://huggingface.co/datasets/textdetox/multilingual_paradetox uses the openrail++ license777The Responsible AI License allows users to take advantage of the model in a wide range of settings (including free use and redistribution) as long as they respect the specific use case restrictions outlined, which correspond to model applications the licensor deems ill-suited for the model or are likely to cause harm..
@@ A.2 Experimental Details
Our experimental setup comprises four learning paradigms across four multilingual LLMs, totaling 504 configurations: (1) zero-shot () evaluation across 9 languages, (2) cross-lingual fine-tuning (-) with 81 language pairs, (3) partial fine-tuning (-) with three data portions per language (27 configurations), and (4) multilingual fine-tuning (-) across 9 languages.

### v2

@@ Abstract
As large language models (LLMs) become increasingly prevalent in global applications, ensuring that they are toxicity-free across diverse linguistic contexts remains a critical challenge. We explore “Cross-lingual Detoxification”, a cross-lingual paradigm that mitigates toxicity, enabling detoxification capabilities to transfer between high and low-resource languages across different script families. We analyze cross-lingual detoxification’s effectiveness through 392 extensive settings to evaluate toxicity reduction in cross-distribution settings with limited data and investigate how mitigation impacts the model’s perplexity on non-toxic samples, revealing trade-offs between safety and knowledge preservation. Our code and dataset are publicly available.111 https://github.com/himanshubeniwal/Breaking-mBad
@@ 1 Introduction
We investigate Cross-Lingual Detoxification (X-DET), a methodology to detoxify language models in a source language and to evaluate transfer effects across seven target languages. We utilize parallel toxic-neutral pairs to perform the detoxification. We showcase this technique that performs efficiently in cross-lingual settings. Our analysis encompasses 392 experimental configurations, comprising 7 languages (49 language pairs), 4 learning strategies, and 4 mLLMs (details in Section A.2).
Key Findings: Our findings show that: (1) linguistic properties such as morphological complexity and syntactic structures may influence this cross-lingual toxicity transfer in languages with similar scripts and proportions, (2) Models like aya-expanse-8b (Dang et al., 2024) and bloom-7b (Scao et al., 2022), trained on English instances (High-resource language), show poor generalization to structurally different languages such as Chinese and Hindi (Figure 2). Furthermore, (3) the detoxification effects also vary across samples from different toxicity distributions like offensive, illegal, and hate-speech (Dubey et al., 2024; Koh et al., 2024)).
@@
Our experiments across 392 configurations show that cross-lingual detoxification significantly outperforms multilingual and proportional fine-tuning approaches.
Cross-distribution detoxification proves effective even with limited parallel data (10%, 20%, and 30% of the entire data), achieving effective detoxification without requiring extensive datasets in similar scripts and pretraining language proportion.
@@ 2 Related Work
Early work on identifying and mitigating toxicity in language models focused primarily on English (Gehman et al., 2020; Xu et al., 2021; Leong et al., 2023; Lee et al., 2024). Initial approaches employed supervised fine-tuning with annotated datasets and keyword-based filtering (Pozzobon et al., 2024; Dementieva et al., 2025b), which often degraded model fluency. While subsequent research introduced preference optimization techniques to align models with safety principles (Li et al., 2024), these studies predominantly target high-resource languages, assuming universal transferability of toxicity pattern (Moskovskiy et al., 2022; Mukherjee et al., 2023; Wang et al., 2024a; Jain et al., 2024; Jiang & Zubiaga, 2024).
@@ 3 Experiments
Problem Setting Let be a set of different languages. Each language is associated with a dataset containing pairs of toxic and non-toxic sentences written in language . Detoxification is the task of using toxic sentences from language to update a language model such that it assigns a low probability to toxic sentences across all languages. More details in
Dataset For our experiments, we utilize the multilingual parallel detoxification dataset: textdetox/multilingual_paradetox222https://huggingface.co/datasets/textdetox/multilingual_paradetox (Bevendorff et al., 2024; Dementieva et al., 2024b; 2025a), which provides parallel toxic and neutral texts across seven333We systematically investigate across the following script families: (1) Latin: German (de), English (en), Spanish (es), (2) Cyrillic: Russian (ru), (3) Devnagri: Hindi (hi), (4) Arabic: Arabic (ar), (5) Han: Chinese (zh). typologically diverse languages. Each language contains carefully curated parallel samples with toxic content paired with its semantically equivalent neutral (Non-toxic) samples. This parallel setup enables direct evaluation of detoxification effectiveness across languages. More details are in Section §A.1.
Models We employ four models to showcase the different behavior and findings: aya-expanse-8B (Dang et al., 2024), aya-23-8B (Aryabumi et al., 2024), mT5-Large (Xue et al., 2021), and bloom-7B1 (Scao et al., 2022). Training details are available in Section §A.2.
@@ 3.1 Learning and Evaluation Paradigms
Cross-lingual Fine-Tuning (-): We finetune each model on 300 neutral-train samples from one source language and evaluate on the test sets across all target languages, yielding 81 language-pair configurations across 7 languages.
@@
Multilingual Fine-Tuning (-): Models are fine-tuned on the aggregated neutral-train samples across all languages (2,700 samples; 300 samples x 7 languages) and evaluated on test sets.
@@ 3.2 Metrics
We prompt the model with the following simple template for all the toxic-train, toxic-test, and neutral-test samples: “Complete the sentence: {SENTENCE}”. We evaluate model performance using two standard metrics: (1) toxicity and (2) perplexity. For toxicity detection, we employ Perspective-API,444https://perspectiveapi.com/ a standardized tool supporting 17 languages across high and low-resource categories, to compute the average toxicity score () across languages (Gehman et al., 2020; Jain et al., 2024). We present the toxicity measurements for the zero-shot () baseline and the corresponding mitigation delta scores555The differential mitigation scores () are calculated by computing the arithmetic difference between the toxicity baseline and the respective fine-tuned variants’ toxicity scores (, where ). () for models fine-tuned with -, -, and -. The model’s perplexity is computed using fine-tuned models. More details are provided in Appendix §A.5.
@@ 4 Results and Discussion
(RQ1) How well does detoxification transfer across languages?
Analysis of the aya-23-8B model in Figure 2 shows superior detoxification in high-resource languages: es (10.41%), de (9.31%), and en (2.34%), with similar trends in aya-expanse-8B (Table 1). Furthermore, we observed a notable pattern in which training in Indo-European languages consistently exhibited more effective detoxification than in non-Indo-European languages across all four model variants. We attribute this disparity to two primary factors: (1) the proportional representation of languages during the pretraining phase, and (2) the inherent similarities in script among related languages. Details in Section §A.3.
@@
(RQ2) Can we effectively mitigate toxicity in cross-lingual settings with limited data?
Figure 3 illustrates the variation in toxicity scores across different training data proportions: , , , and (-), where we finetune on the portion of languages and report the AT over a specific language. Notably, ar demonstrated improved detoxification performance, aligning with the trends observed in en and es. Our analysis of these languages’ behavior, presented in Figures 7 and 8 (detailed further in Section §A.4), reveals that the fine-tuning causes the embedding representations to converge, suggesting increased similarity in the model’s handling of toxicity across these languages.
@@
(RQ3) How does cross-lingual detoxification impact perplexity? Our perplexity analysis reveals that Indo-European languages, particularly hi, show improved scores (9.01) in aya-expanse-8B’s toxic-train split (Table 15), though both - and - negatively impacted overall perplexity across models (More details in Section §A.5). Embedding similarity analysis before and after detoxification indicates a shift in the relationship between en and de, with their similarity score decreasing to 0.69 in Figures 7 and 8.
@@ Acknowledgments
Our work explores the challenges of Large Language Models (LLMs) in generating toxic content across different language families, including Indo-European, Non-Indo-European, and Right-to-Left script languages. Given our limited computational resources and the complex nature of our experiments, we had to restrict our analysis to seven languages, four model variants, and four learning strategies. Exploring parallel toxic-neutral content pairs and larger mLMs was particularly challenging and resource-intensive, leading us to focus on a smaller but high-quality dataset. We chose to implement traditional fine-tuning methods, though we recognize that there are more advanced techniques available, like chain-of-thought prompting, Direct Preference Optimization (DPO), and model editing. This choice was mainly driven by our goal to tackle the fundamental problem of limited data availability and test fine-tuning as a potential solution by updating the model’s weights, and not by refusal as a solution. Furthermore, the models are susceptible to jailbreaking, adversarial attacks, and using toxic refusal (ex., “Sorry I cannot respond..”) (Morris et al., 2020). Thus, we prioritized weight updation as a strategy. Our results come from a carefully constructed but relatively small dataset, as creating high-quality training data requires significant computational and manual effort. Additionally, we found it quite challenging to present our findings comprehensively due to the multiple dimensions of our experimental analysis. Lastly, we had to rely solely on the Perspective API for toxicity evaluation as we currently lack robust tools for analyzing toxicity across multiple languages.
@@ A.1 Dataset Split
From the original set, we create our experimental splits by sampling 400 pairs, constructing a training set of 300 parallel pairs (toxic-train and neutral-train) and a test set of 100 pairs (toxic-test and neutral-test). We utilize the 300 neutral-train pairs to fine-tune and evaluate our hypothesis of cross-lingual detoxification using straightforward neutral samples. Further, the textdetox/multilingual_paradetox dataset666https://huggingface.co/datasets/textdetox/multilingual_paradetox uses the openrail++ license777The Responsible AI License allows users to take advantage of the model in a wide range of settings (including free use and redistribution) as long as they respect the specific use case restrictions outlined, which correspond to model applications the licensor deems ill-suited for the model or are likely to cause harm..
@@ A.2 Experimental Details
Our experimental setup comprises four learning paradigms across four multilingual LLMs, totaling 392 configurations: (1) zero-shot () evaluation across 7 languages, (2) cross-lingual fine-tuning (-) with 81 language pairs, (3) partial fine-tuning (-) with three data portions per language (27 configurations), and (4) multilingual fine-tuning (-) across 7 languages.